OpenStudy (anonymous):

An inflated beach ball with radius r = 25 centimeters holds approximately 65,400 cubic centimeters of air. When the ball is compressed, the radius is 2/3 the original size. Approximately how much air does the partially deflated ball now hold?

OpenStudy (anonymous):

19,400 cm^3 14,500 cm^3 1,200 cm^3 4,800 cm^3

OpenStudy (anonymous):

@SweetsAndEverything @electrokid

OpenStudy (anonymous):

we do not need the exact formula.. just understand that volume is proportional to the cube of the radius \[ V\propto r^3\\ {V_1\over V_2}={r_1^3\over r_2^3} \] V1 = 65400, r1=25 V2=? r2=2*25/3